# MDD-5k: A New Diagnostic Conversation Dataset for Mental Disorders Synthesized via Neuro-Symbolic LLM Agents

### v1

@@ Abstract
The clinical diagnosis of most mental disorders primarily relies on the conversations between psychiatrist and patient. The creation of such diagnostic conversation datasets is promising to boost the AI mental healthcare community. However, directly collecting the conversations in real diagnosis scenarios is near impossible due to stringent privacy and ethical considerations. To address this issue, we seek to synthesize diagnostic conversation by exploiting anonymous patient cases that are easier to access. Specifically, we design a neuro-symbolic multi-agent framework for synthesizing the diagnostic conversation of mental disorders with large language models. It takes patient case as input and is capable of generating multiple diverse conversations with one single patient case. The framework basically involves the interaction between a doctor agent and a patient agent, and achieves text generation under symbolic control via a dynamic diagnosis tree from a tool agent. By applying the proposed framework, we develop the largest Chinese mental disorders diagnosis dataset MDD-5k, which is built upon 1000 cleaned real patient cases by cooperating with a pioneering psychiatric hospital, and contains 5000 high-quality long conversations with diagnosis results as labels. To the best of our knowledge, it’s also the first labelled Chinese mental disorders diagnosis dataset. Human evaluation demonstrates the proposed MDD-5k dataset successfully simulates human-like diagnostic process of mental disorders. The dataset and code will become publicly accessible in https://github.com/lemonsis/MDD-5k.
22footnotetext: Work done during an internship.33footnotetext: Corresponding authors.
@@ Introduction
Mental health issues have garnered increasing attention in recent years. According to the statistics of World Health Organization (WHO), one in every eight people in the world lived with a mental disorder in 2019, and people living with anxiety and depressive disorders kept rising significantly because of the COVID-19 pandemic (Organization et al. 2022). With the recent progress of large language models (LLMs) (Ouyang et al. 2022; OpenAI 2024), which emerges capabilities of relevant context incorporation and human-like text generation, many researchers turn to building conversational AI system for mental healthcare. Current implementations can be divided into two categories, finetuning a small model (e.g. Llama2-7B (Touvron et al. 2023)) with physician-patient conversations (EmoLLM 2024; Yang et al. 2024b; Liu et al. 2023) or building prompt-based physician-patient role-playing framework (Wang et al. 2024a; Zhang et al. 2024) with state-of-the-art language models (e.g. ChatGPT (Ouyang et al. 2022)). Regardless of the method employed, domain-specific mental health datasets play a fundamental and indispensable role.
@@
We focus on diagnostic conversation dataset for mental disorders in this work. The clinical diagnosis of mental disorders differs from other diseases in that it primarily relies on the mental status examination of patients, which is reflected through conversations between psychiatrists and patients rather than physiological indices (First 2014). Therefore, the collection of mental disorders diagnostic conversation is promising to facilitate a variety of downstream tasks in AI mental health research like auxiliary diagnosis chatbot, mental disorders classification, etc. However, while many previous studies (Sun et al. 2021; Chen et al. 2023b) focused on emotional support or psychological counseling data, few work shed light on diagnostic conversations of mental disorders. This can be attributed to two main factors. First, diagnostic conversations in real scenarios are extremely hard to acquire due to the privacy and ethical consideration. Second, synthesizing diagnostic conversations from scratch is also challenging. Unlike psychological counseling or empathetic dialogue, diagnosis follows standardized process and requires professional medical knowledge. Consequently, directly employing LLMs for data synthesis in this context often yields poor outcomes (Tu et al. 2024). $\text{D}^{4}$ (Yao et al. 2022) made the first attempt by simulating diagnostic conversations with employed workers. However, it only covers depressive disorder and entirely depends on human annotation. The generated content is also short and far from similar to diagnostic conversations in real scenarios.
@@
We propose a neuro-symbolic multi-agent framework that takes patient cases as input to synthesize diagnostic conversations of mental disorders. The framework involves three types of large language model agents: a doctor agent, a patient agent, and a symbolic tool agent responsible for managing diagnostic topic shift. This framework features two major innovations: (i) One-to-many patientcase-to-dialogue generation that maximizes the utilization of precious real patient cases. As shown in Figure 1, unlike previous studies (Zhang et al. 2024; Wang et al. 2024a) that generate one conversation with one patient case. Our proposed framework is capable of generating multiple diverse diagnostic conversations with one single patient case. Specifically, three methods ensure the diversity and correctness of diagnostic process. First, doctor agents with different diagnosis habits are designed and randomly selected for each conversation. Second, we use LLM with knowledge graph to generate multiple fictitious patient experiences based on one patient case. The patient experiences serve as background information for patient agents during generation. Since the diagnosis of mental disorders mainly relies on symptoms rather than concrete events, integrating the fictitious patient experiences enhances the diversity of synthesized conversation while maintaining the accuracy of the diagnostic process. Third, the sequence of diagnostic topics is randomly determined for each conversation. (ii) Another significant innovation lies in text generation under symbolic control via a dynamic diagnosis tree. This tree consists of a fixed symptom inquiry tree and a dynamic experience inquiry tree. Clinical diagnosis of mental disorders strictly follows standards from ICD-11 (Organization et al. 2018) or DSM-5 (American Psychiatric Association et al. 2013). To simulate this process, we design the fixed symptom inquiry tree based on Structured Clinical Interview for DSM-5 (SCID-5) (First 2014), covering all the diagnostic topics for important symptoms inquiry. The experience inquiry tree is constructed by extracting possible topics from patient’s response of past experiences. It’s designed to establish deeper engagements with the patient.
@@
By applying the proposed framework, we release the largest Chinese Mental Disorder Diagnosis dataset MDD-5k. It’s also the first labelled mental disorders diagnostic conversation dataset with diagnosis results from professional psychiatrists. MDD-5k contains 5000 high-quality diagnostic conversations, with an averaged 26.8 turns and 6906.8 Chinese words per dialogue. It is built upon 1000 real patient cases from a pioneering psychiatric hospital, covering more than 25 different diseases. All the patient cases have been cleaned and filtered in accordance with global standards to ensure the complete protection of patient private information. The MDD-5k dataset will become publicly accessible when the ethical review finishes.
@@
  • The largest Chinese mental disorders diagnosis dataset MDD-5k is proposed, which contains 5000 high-quality long conversations with convincing diagnosis results as labels. To the best of our knowledge, it’s also the first mental disorders diagnosis dataset with labels.
@@
The synthesis process of mental disorders diagnostic conversations is presented. The neuro-symbolic multi-agent framework features one-to-many patientcase-to-dialogue generation, indicating multiple generated conversations are based on the same patient case. As shown in Figure 2, the framework basically involves the interactions among a doctor agent, a patient agent, and a tool agent. All the agents are played by large language models (LLMs). The doctor agent controlled by a dynamic diagnosis tree leads the diagnostic topic shift of the whole conversation. The patient agent responds to the doctor agent based on the preprocessed patient case and fictitious patient experience generated by the tool agent. The tool agent is also responsible for the several symbolic operations of dynamic diagnosis tree.
@@
The quality of patient cases is vital to the diagnostic conversation synthesis. We cooperate with a pioneering psychiatric hospital and obtain over 1000 real cases of patients with mental disorders. All these patient cases have undergone data masking to prevent the leakage of sensitive personal information. The data masking process follows the standards below: (i) Private information of patients (e.g. name, date of birth, date of examination, etc.) is removed from the patient case. (ii) Patient age is rounded to the nearest ten. For example, the age of a 24-year-old patient is 20 on the preprocessed patient case. (iii) All the concrete locations are replaced with vague ones. The above preprocessing steps strictly follow the Chinese information security technology guide for health data security (GB/T 39725-2020). A complete cleaned patient case is illustrated in Appendix.
@@
We perform one-to-many patientcase-to-dialogue generation, which indicates one patient case will be applied to generate multiple diagnostic conversations. One key factor contributes to the diversity of generated conversations is the patient experience. It specifically refers to the past experiences that directly or indirectly lead to the mental illness problem of patients in this paper. The diagnosis of mental disorders differs from other illnesses in that it mainly depends on the conversations between psychiatrists and patients instead of physiological indices. Psychiatrists provide diagnosis result and treatment based on the acquired symptoms of patients during communicating. As a result, if the correspondence between symptoms and diagnosis can be assured, the correctness and quality of the synthesized diagnostic conversation is guaranteed and is not affected by detailed patient experience. In this sense, it’s feasible to generate multiple patient experiences with one patient case for synthesizing multiple diagnostic conversations.
Figure 2: The neuro-symbolic multi-agent framework for synthesizing diagnostic conversation of mental disorders.
@@
The neuro-symbolic dynamic diagnosis tree is managed by the tool agent and offers operations for guiding both the doctor agent and the patient agent. Some operations are implemented by LLM and some are by rules. We first define five data types: Text, LNode, Tree, Graph, Bool. Text refers to natural language text. LNode stands for the leaf node of a dynamic diagnosis tree, indicating diagnostic topic in conversation. Tree refers to the hierarchical tree structure with a set of connected nodes. Although LNode can be viewed as a special Tree or Text, we treat it as a separate data type for clearer expression. Graph specifically refers to the knowledge graph for fictitious patient experience generation as explained before. Bool is a boolean variable is either true or false. The operations for a doctor agent include:
@@
Table 1: Statistics of different datasets. Avg. words #dial measures the average Chinese words per dialogue. Avg. words #doc and Avg. words #pat measure the average Chinese words per doctor’s response and patient’s response.
The doctor agent generates responses through operation DocGen($ln$: LNode, $t^{(1)}$: Text) $\rightarrow$ $t^{(2)}$: Text, which takes current diagnostic topic $ln$ and dialogue history $t^{(1)}$ as input and outputs response $t^{(2)}$. To further improve the diversity of generated conversations, we design different diagnosis habits for the doctor agent. The diagnostic habits contain age, gender, specialties, empathetic dialogue, diagnosis speed, explanation, and serve as persona prompt for the doctor agent. An example is shown in Figure 2. Specifically, the factors of empathetic dialogue and diagnosis speed exert huger effect on the doctor’s response. If the doctor agent is accustomed to communicate empathetically, the EmpathGen operation introduced before will replace the DocGen operation for generation. If the diagnosis speed is set as fast, the doctor agent will speed up the diagnosis process, which leads to shorter conversations.
Figure 3: Patient information of the MDD-5 dataset.
As to the patient agent, since the doctor agent leads the diagnosis process, the patient agent is designed to passively respond to the doctor based on known knowledge, including the patient case and generated fictitious experience. The operation PatGen($ln$: LNode, $t^{(1)}$: Text, $t^{(2)}$: Text) $\rightarrow$ $t^{(3)}$: Text is responsible for patient response generation, which takes current diagnostic topic $ln$, dialogue history $t^{(1)}$, and patient case information $t^{(2)}$ as input and outputs proper response $t^{(3)}$. If the patient agent determines to respond with personal experience under the control of dynamic diagnosis tree, fictitious patient experience will be fused into the patient case as input $t^{(2)}$. The whole process of the multi-agent framework for diagnostic conversation simulation of mental disorders is detailedly shown through pseudo-code in Algorithm 1 in the Appendix.
@@
The MDD-5k dataset is generated through the neuro-symbolic multi-agent framework with 1000 real patient cases. 5 different fictitious patient experiences are generated by OpenAI gpt-4o (OpenAI 2024) based on 1 patient case, which leads to a total of 5000 experiences corresponding to the 5000 conversations in the dataset. We also create 5 doctors with different diagnosis habits, and 1 doctor will be randomly picked for generating each conversation in the dataset. Since the patient cases are still under the ethical review of psychiatric hospital, we randomly select 20 available patient cases and generate 100 conversations by gpt-4o for evaluation. The other 4900 conversations are currently generated by Qwen2-72B-Instruct (Yang et al. 2024a) deployed on NVIDIA A100-80G GPUs locally. We will apply gpt-4o for generating the rest 4900 conversations when the ethical review finishes.
@@
Five annotators are employed for the human evaluation. Three of the them are professional psychiatrists who have years of clinical experience. The other two annotators are experienced in mental health data processing. The evaluation is conducted in a double-blind manner.
@@
The statistical details of MDD-5k and other compared datasets are presented in Table 1. The MDD-5k dataset contains diagnostic conversations covering over 25 mental health illnesses. It includes 5000 dialogues, each comprising an averaged of 6906.8 Chinese words which is almost ten times the compared datasets. The average dialogue turns are 26.8, slightly longer than the 21.6 turns of $\text{D}^{4}$. It is also a labelled dataset with diagnosis result from professional psychiatrist as label for each conversation. Compared to the direct role-playing method without applying the multi-agent framework, the generated doctor response is about the same length. But the dialogue turns and patient response of MDD-5k are significantly longer, highlighting the effectiveness of our proposed framework in diagnostic conversation simulation. In the case study presented in Appendix, we show three complete samples of conversation with corresponding doctor persona, patient case, fictitious patient experience and the dynamic diagnosis tree.
@@
We design a neuro-symbolic multi-agent framework for synthesizing diagnostic conversation of mental disorders, and apply it for building the first and largest open-source Chinese mental disorders diagnosis dataset with diagnosis results as labels. The framework features controllable one-to-many patientcase-to-dialogue generation. Conversation between a doctor agent and a patient agent is guided by a dynamic diagnosis tree. We also apply several techniques to improve the diversity of generated conversations. Human evaluation shows the quality of the proposed MDD-5k dataset exceeds existing relevant datasets on seven indicators. The MDD-5k dataset is believed to contribute to a wide range of downstream tasks like mental disorders classification, mental disorders diagnosis assistant training, etc.
@@
Ethics Statement
The collection of patient cases was conducted at the Shanghai Mental Health Center. All patients were informed that their information would be collected and used exclusively for research purposes. As detailed in the Patient Cases Preprocessing section, all data masking procedures strictly follow the Chinese information security technology guidelines for health data security (GB/T 39725-2020). Currently, the patient cases and the synthesized diagnostic conversation dataset, MDD-5k, are undergoing an ethics review at the Shanghai Mental Health Center. We plan to release the MDD-5k dataset for research purposes only after the ethics review finishes. To prevent any potential privacy data leakage, all experiments are conducted on servers located within the Shanghai Mental Health Center.

### v2

@@ Abstract
The clinical diagnosis of most mental disorders primarily relies on the conversations between psychiatrist and patient. The creation of such diagnostic conversation datasets is promising to boost the AI mental healthcare community. However, directly collecting the conversations in real diagnosis scenarios is near impossible due to stringent privacy and ethical considerations. To address this issue, we seek to synthesize diagnostic conversation by exploiting anonymized patient cases that are easier to access. Specifically, we design a neuro-symbolic multi-agent framework for synthesizing the diagnostic conversation of mental disorders with large language models. It takes patient case as input and is capable of generating multiple diverse conversations with one single patient case. The framework basically involves the interaction between a doctor agent and a patient agent, and generates conversations under symbolic control via a dynamic diagnosis tree. By applying the proposed framework, we develop the largest Chinese mental disorders diagnosis dataset MDD-5k. This dataset is built upon 1000 real, anonymized patient cases by cooperating with Shanghai Mental Health Center and comprises 5000 high-quality long conversations with diagnosis results and treatment opinions as labels. To the best of our knowledge, it’s also the first labeled dataset for Chinese mental disorders diagnosis. Human evaluation demonstrates the proposed MDD-5k dataset successfully simulates human-like diagnostic process of mental disorders.
Code&Datasethttps://github.com/lemonsis/MDD-5k.
@@ Introduction
Mental health issues have garnered increasing attention in recent years. According to the statistics of World Health Organization (WHO), one in every eight people in the world lived with a mental disorder in 2019, and people living with anxiety and depressive disorders kept rising significantly because of the COVID-19 pandemic (WHO et al. 2022). With the recent progress of large language models (LLMs) (Ouyang et al. 2022), which emerges capabilities of human-like text generation, many researchers turn to building conversational AI system for mental healthcare. Current implementations can be divided into two categories, finetuning a small model (e.g. Llama2-7B (Touvron et al. 2023)) with physician-patient conversations (EmoLLM 2024; Yang et al. 2024b; Liu et al. 2023) or building prompt-based physician-patient role-playing framework (Wang et al. 2024a; Zhang et al. 2024) with state-of-the-art language models (e.g. ChatGPT (Ouyang et al. 2022)). Regardless of the method employed, domain-specific mental health datasets play a fundamental and indispensable role.
@@
We focus on diagnostic conversation dataset for mental disorders in this work. The clinical diagnosis of mental disorders differs from other diseases in that it primarily relies on the mental status examination of patients, which is reflected through conversations between psychiatrists and patients rather than physiological indices (First 2014). Therefore, the collection of mental disorders diagnostic conversation is promising to facilitate a variety of downstream tasks in AI mental health research like auxiliary diagnosis chatbot, mental disorders classification, etc. However, while many previous studies (Sun et al. 2021; Chen et al. 2023b) focused on emotional support or psychological counseling data, few work shed light on diagnostic conversations of mental disorders. This can be attributed to two main factors. First, diagnostic conversations in real scenarios are extremely hard to acquire due to the privacy and ethical consideration. Second, synthesizing diagnostic conversations from scratch is also challenging. Unlike psychological counseling or empathetic dialogue, diagnosis follows standardized process and requires professional medical knowledge. Consequently, directly employing LLMs for data synthesis in this context often yields poor outcomes (Tu et al. 2024). $\text{D}^{4}$ (Yao et al. 2022) made the first attempt by simulating diagnostic conversations with employed annotators. However, it only covers depressive disorder and entirely depends on human annotation. The generated content is also short and far from similar to diagnostic conversations in real scenarios.
@@
We propose a neuro-symbolic multi-agent framework that takes patient cases as input to synthesize diagnostic conversations of mental disorders. The framework involves three types of large language model agents: a doctor agent, a patient agent, and a symbolic tool agent responsible for managing diagnostic topic shift. This framework features two major innovations: (i) One-to-many patientcase-to-dialogue generation that maximizes the utilization of precious real patient cases. As shown in Figure 1, unlike previous studies (Zhang et al. 2024; Wang et al. 2024a) that generate one conversation with one patient case. Our proposed framework is capable of generating multiple diverse diagnostic conversations with one single patient case. Specifically, three methods ensure the diversity and correctness of diagnostic process. First, doctor agents with different diagnosis habits are designed and randomly selected for each conversation. Second, we use LLM with knowledge graph to generate multiple fictitious patient experiences given one patient case. The patient experiences serve as background information for patient agents during generation. Since the diagnosis of mental disorders mainly relies on symptoms rather than concrete events, integrating the fictitious patient experiences enhances the diversity of synthesized conversation while maintaining the accuracy of the diagnostic process. Third, the sequence of diagnostic topics is randomly determined for each conversation. (ii) Another significant innovation lies in text generation under symbolic control via a dynamic diagnosis tree. This tree consists of a fixed symptom inquiry tree and a dynamic experience inquiry tree. Clinical diagnosis of mental disorders strictly follows standards from ICD-11 (Organization et al. 2018) or DSM-5 (American Psychiatric Association et al. 2013). To simulate this process, we design a fixed symptom inquiry tree based on Structured Clinical Interview for DSM-5 (SCID-5) (First 2014), covering all the diagnostic topics for important symptoms inquiry. The experience inquiry tree is constructed by extracting possible topics from patient’s response of past experiences. It’s designed to establish deeper engagements with the patient.
@@
By applying the proposed framework, we release the largest Chinese Mental Disorder Diagnosis dataset MDD-5k. It’s also the first labeled mental disorders diagnostic conversation dataset with diagnosis results from professional psychiatrists. MDD-5k contains 5000 high-quality diagnostic conversations and is built upon 1000 anonymized, real patient cases from Shanghai Mental Health Center, covering over 25 different diseases. All the patient cases are cleaned and filtered in accordance with global standards to ensure the complete protection of private information.
@@
  • We propose MDD-5k which is the largest and first labeled Chinese mental disorders diagnosis dataset to the best of our knowledge.
@@
The synthesis process of mental disorders diagnostic conversations is presented. As shown in Figure 2, the framework basically involves the interactions among a doctor agent, a patient agent, and a tool agent. All the agents are played by large language models (LLMs). The doctor agent controlled by a dynamic diagnosis tree leads the diagnostic topic shift of the whole conversation. The patient agent responds to the doctor agent based on the preprocessed patient case and fictitious patient experience generated by the tool agent. The tool agent is also responsible for several symbolic operations of the dynamic diagnosis tree.
@@
The quality of patient cases is vital to the diagnostic conversation synthesis. We cooperate with Shanghai Mental Health Center and obtain over 1000 real cases of patients with mental disorders. All these patient cases have undergone data masking to prevent the leakage of sensitive personal information. The data masking process follows the standards below: (i) Private information of patients (e.g. name, date of birth, date of examination, etc.) is removed from the patient case. (ii) Patient age is rounded to the nearest ten. For example, the age of a 24-year-old patient is 20 on the preprocessed patient case. (iii) All the concrete locations are replaced with vague or fake ones. The above preprocessing steps strictly follow the Chinese information security technology guide for health data security (GB/T 39725-2020).
@@
We perform one-to-many patientcase-to-dialogue generation, which indicates one patient case will be applied to generate multiple diagnostic conversations. One key factor contributes to the diversity of generated conversations is the patient experience. It specifically refers to the past experiences that directly or indirectly lead to the mental illness problem of patients in this paper. The diagnosis of mental disorders differs from other illnesses in that it mainly depends on the conversations between psychiatrists and patients instead of physiological indices (First 2014). Psychiatrists provide diagnosis result and treatment based on the acquired symptoms of patients during communicating. As a result, if the correspondence between symptoms and diagnosis can be assured, the correctness and quality of the synthesized diagnostic conversation is guaranteed and is not affected by detailed patient experience. In this sense, it’s feasible to generate multiple patient experiences with one patient case for synthesizing multiple diagnostic conversations.
Figure 2: The neuro-symbolic multi-agent LLM framework for synthesizing diagnostic conversation of mental disorders.
@@
The neuro-symbolic dynamic diagnosis tree is managed by the tool agent and offers operations for guiding both the doctor agent and the patient agent. Some operations are implemented by LLM and some are by rules. We first define five data types: Text, LNode, Tree, Graph, Bool. Text refers to natural language text. LNode stands for the leaf node of a dynamic diagnosis tree, indicating diagnostic topic in conversation. Tree refers to the hierarchical tree structure with a set of connected nodes. Although LNode can be viewed as a special Tree or Text, we treat it as a separate data type for clearer expression. Graph specifically refers to the knowledge graph for fictitious patient experience generation as explained before. Bool is a boolean variable which is either true or false. The operations for a doctor agent include:
@@
Table 1: Statistics of different datasets. Avg. words #dial measures the average Chinese characters per dialogue. Avg. words #doc and Avg. words #pat measure the average Chinese characters per doctor’s response and patient’s response.
The doctor agent generates responses through operation DocGen($ln$: LNode, $t^{(1)}$: Text) $\rightarrow$ $t^{(2)}$: Text, which takes current diagnostic topic $ln$ and dialogue history $t^{(1)}$ as input and outputs response $t^{(2)}$. To further improve the diversity of generated conversations, we design different diagnosis habits for the doctor agent. The diagnosis habits contain age, gender, specialties, empathetic dialogue, diagnosis speed, explanation, and serve as persona prompt for the doctor agent. An example is shown in Figure 2. Specifically, the factors of empathetic dialogue and diagnosis speed exert huger effect on the doctor’s response. If the doctor agent is accustomed to communicate empathetically, the EmpathGen operation introduced before will replace the DocGen operation for generation. If the diagnosis speed is set as fast, the doctor agent will speed up the diagnosis process, which leads to shorter conversations.
Figure 3: Patient information of the MDD-5k dataset.
As to the patient agent, since the doctor agent leads the diagnosis process, the patient agent is designed to passively respond to the doctor based on known knowledge, including the patient case and generated fictitious experience. The operation PatGen($ln$: LNode, $t^{(1)}$: Text, $t^{(2)}$: Text) $\rightarrow$ $t^{(3)}$: Text is responsible for patient response generation, which takes current diagnostic topic $ln$, dialogue history $t^{(1)}$, and patient case information $t^{(2)}$ as input and outputs proper response $t^{(3)}$. If the patient agent determines to respond with personal experience under the control of dynamic diagnosis tree, fictitious patient experience will be fused into the patient case as input $t^{(2)}$. The whole process of the multi-agent framework for diagnostic conversation simulation of mental disorders is detailedly shown through pseudo-code in the Appendix.
@@
The MDD-5k dataset is generated through the neuro-symbolic multi-agent framework with 1000 real patient cases. 5 different fictitious patient experiences are generated by gpt-4o based on 1 patient case, which leads to a total of 5000 experiences corresponding to the 5000 conversations in the dataset. We also create 5 doctors with different diagnosis habits, and 1 doctor will be randomly picked for generating each conversation in the dataset. Since the patient cases are still under the ethical review, we randomly select 20 available patient cases and generate 100 conversations with gpt-4o for evaluation. The other 4900 conversations are currently generated by Qwen2-72B-Instruct (Yang et al. 2024a) deployed on NVIDIA A100-80G GPUs locally.
@@
Five annotators are employed for the human evaluation. Three of the them are professional psychiatrists who have years of clinical experience. The other two annotators are experienced in mental health data processing. The evaluation is conducted in a blind manner.
@@
The statistical details of MDD-5k and other compared datasets are presented in Table 1. The MDD-5k dataset contains diagnostic conversations covering over 25 mental health illnesses. It includes 5000 dialogues, each comprising an averaged of 6906.8 Chinese characters which is almost ten times to compared datasets. The average dialogue turns are 26.8, slightly longer than the 21.6 turns of $\text{D}^{4}$. MDD-5k is also a labeled dataset with diagnosis result and treatment opinion from professional psychiatrist as label for each conversation, while $\text{D}^{4}$ only contains diagnosis results. Compared to the direct role-playing method without applying the multi-agent framework, the generated doctor response is about the same length. But the dialogue turns and patient response of MDD-5k are significantly longer, highlighting the effectiveness of our proposed framework in diagnostic conversation simulation. In the case study presented in Appendix, we show three complete samples of conversation with corresponding doctor persona, patient case, fictitious patient experience and the dynamic diagnosis tree.
@@
We design a neuro-symbolic multi-agent framework for synthesizing diagnostic conversation of mental disorders, and apply it for building the first and largest open-source Chinese mental disorders diagnosis dataset with diagnosis results and treatment opinions as labels. The framework features controllable one-to-many patientcase-to-dialogue generation. Conversation between a doctor agent and a patient agent is guided by a dynamic diagnosis tree. We also employ several techniques to improve the diversity of generated conversations. Human evaluation shows the quality of the proposed MDD-5k dataset exceeds compared datasets on seven indicators. The MDD-5k dataset is believed to contribute to a wide range of downstream tasks like mental disorders classification, mental disorders diagnosis assistant training, etc.
@@
Ethical Statement
The collection of patient cases was conducted at the Shanghai Mental Health Center. All patients were informed that their information would be collected and used exclusively for research purposes. As detailed in the Patient Cases Preprocessing section, all data masking procedures strictly follow the Chinese information security technology guidelines for health data security (GB/T 39725-2020). Currently, the patient cases and the synthesized diagnostic conversation dataset, MDD-5k, are undergoing an ethics review. We plan to release the MDD-5k dataset for research purposes only after the ethics review finishes. To prevent any potential privacy data leakage, all experiments are conducted on servers located within the Shanghai Mental Health Center.